OpenStudy (anonymous):

Solve the following initial value problem dy/dt = 1-y^2 y(0)=1

OpenStudy (anonymous):

@Loser66 @surjithayer @ganeshie8

OpenStudy (anonymous):

(In case you didn't notice, I'm not too sure how to solve the integral 1/(1-y^2) )

OpenStudy (anonymous):

(If I were to use separation of variables..)

OpenStudy (anonymous):

(And I can't do linear case either cause y^2 is obviously not linear)..

hartnn (hartnn):

partial fractions

hartnn (hartnn):

1/(1-y)(y+1) = A/(1-y) + B/(1+y) find A,B

OpenStudy (anonymous):

\[\frac{ dy }{1-y^2 }=dt, \frac{ 1 }{1-y^2 }=\frac{ 1 }{ \left( 1-y \right)\left( 1+y \right) }=\frac{ 1 }{ \left( 1-y \right)\left( 1+1 \right) }+\frac{ 1 }{\left(1-(-1) \right)\left( 1+y \right) }\] \[=\frac{ 1 }{2\left( 1-y \right) }+\frac{ 1 }{2\left( 1+y \right) }\] \[\frac{ 1 }{ 2 }\left( \frac{ 1 }{ 1-y } +\frac{ 1 }{ 1+y }\right)dy=dt\] integrating both sides \[-\ln \left( 1-y \right)+\ln \left( 1+y \right)=2t+c\] \[\ln \frac{ 1+y }{ 1-y }=2t+c\] when t=0,y=1 here is problem for 1-1=0

OpenStudy (anonymous):

\[1+y=\left( 1-y \right) e ^{2t+c}=C \left( 1-y \right)e ^{2t}\] C can not be found by these values.